OpenStudy (anonymous):

Differentiate y=((x+1)^3)/x^(3/2)

OpenStudy (loser66):

open (x+1)^3 = x^3 + 3x^2 +3x+1 then time x^(-3/2) to the polynomial , then take derivative term by term. No quotient rule. Or if you good at quotient rule, just apply.

OpenStudy (anonymous):

can u do it using quotient rule...m kinda stuck there!!

OpenStudy (loser66):

what is the quotient rule ?

OpenStudy (anonymous):

assume u=(x+1)^3 v=x^3/2 then use the formula.y'={(u').v-(v').u}/v^2

OpenStudy (loser66):

ok, so , what is u'?

OpenStudy (anonymous):

derivative of u

OpenStudy (anonymous):

m attaching the ans

OpenStudy (anonymous):

wich incidentally...m not getin

OpenStudy (loser66):

hopefully I can check! jut post

OpenStudy (anonymous):

posted

OpenStudy (loser66):

that's the answer?

OpenStudy (anonymous):

yup

OpenStudy (loser66):

why don't you post the whole work?

OpenStudy (anonymous):

there's a question.....then this answer....no steps ...no anything!!

OpenStudy (loser66):

why does it mean? it means we have to do it by ourselves, right?

OpenStudy (anonymous):

yup....and u want me 2 do it??....and I want sum1 to help me get there!!

OpenStudy (loser66):

I don't want to do it, it's hard to me. But if you do, I will follow and help to check your steps. Is it ok? If we cannot get that answer, then we tag other for help. is it ok?

OpenStudy (anonymous):

a'ight

OpenStudy (anonymous):

here goes

OpenStudy (loser66):

wut wut... can you calculate u'=?

OpenStudy (anonymous):

calculatin

OpenStudy (anonymous):

vl post the final ans....u just hav to bring it to the ans form,OK???

OpenStudy (loser66):

I think we have to calculate u', v' from the form of rule \[\huge \dfrac{u'v-v'u}{v^2}\]where \( u= (x+1)^3\) and \(v = x^{\frac{3}{2}}\)

OpenStudy (loser66):

after calculating, we plug the results into that form to get the answer

OpenStudy (anonymous):

xactly

OpenStudy (loser66):

so, you should do it, right? I don't know how to do, hihihi

OpenStudy (anonymous):

forgot to write one thing

OpenStudy (anonymous):

the numerator 3x/2 where x shud have the xponential power 1/2 ....add that

OpenStudy (loser66):

the v' at the second term of numerator seems not right, friend. (x^(3/2))' = 3/2 x^(3/2-1)= \(\frac{3}{2} x ^{1/2}\)

OpenStudy (anonymous):

I just corrected it...in the previous comment

OpenStudy (loser66):

so, your problem is not how to simplify it? and the denominator is x^3, right?

OpenStudy (anonymous):

yup how to get x^5/2 in the denom??

OpenStudy (anonymous):

tnx man got...it

OpenStudy (loser66):

I really don't know, friend.

OpenStudy (anonymous):

no probs

OpenStudy (loser66):

so sorry.

OpenStudy (anonymous):

had to observe a bit....tnx for forcing me to stay on it!!......got the ANSWER

OpenStudy (loser66):

when you have a bad helper like me, you have to do everything by yourself, hihihi..

OpenStudy (anonymous):

I am now posting the solution

OpenStudy (loser66):

no need to do that. I am not curious on it. keep it for you only.

OpenStudy (anonymous):

no prob....A push was wat I neede...m a bit lethargic..